A spring with a constant of 200 N/m stretches by 0.03m. What is the potential energy of the spring


1
Expert's answer
2022-05-24T13:56:06-0400
"E_p=\frac{kx^2}{2}\\
=\frac{200*0.03^2}{2}=0.09\:\rm J"